Solve each equation or inequality. Round to four decimal places if necessary.
1. 272p+1= 3​4p-1
Answer #1

272p+1= 3​4p-1

Converting  272p+1 to base 3

(33)2p+1 = 3​4p-1

Now,

33(2p+1) = 3​4p-1

We know that,

if af(x) = ag(x) then f(x) = g(x)

Therefore,

3(2p+1) = 4p-1

6p + 3 = 4p - 1

6p - 4p = -1 - 3

2p = -4

p = -2 Answer

Hope This Helps !